# Mirror Descent Policy Optimisation for Robust Constrained Markov Decision Processes

### v1

@@ 5.1 Cartpole
[…]
@@
where $\mu(s,a)$ is the deterministic next state given $(s,a)\in\mathcal{S}\times\mathcal{A}$ based on the mechanics of the original Cartpole problem. We first train and evaluate algorithms in a CMDP problem for their ability to adhere to safety constraints. We then also show the benefits of robustness training and evaluate algorithms in an RCMDP problem.
@@
To incorporate robustness into our setting, we introduce delta-perturbations as in Wang and Petrik, (2024), using an uncertainty set with parametrisation
@@
where $\delta\in[-\kappa_{i},\kappa_{i}]_{i=1}^{n}$. Our parameter settings differ in that we use a larger uncertainty set, with $\kappa_{i}$ being increased fivefold (making the problem more challenging) while the number of time steps is reduced to make the constraint satisfiable. The robust algorithms are trained on this uncertainty set, where the transition dynamics are adjusted based on the transition mirror ascent algorithm of Wang and Petrik, (2024), i.e. by using Monte Carlo and mirror ascent. In the experiment, this amounts to a projected update, where the projection restricts the update to lie within the uncertainty set.
@@
The training performance of non-robust algorithms can be found in Appendix E.1 (Figure 7). The training performance of robust methods, which use LTMA, can be found in Figure 8. A trade-off in reward vs constraint-cost can be observed; that is, the constrained algorithms put emphasis on reducing the constraint-cost while the unconstrained only maximise the reward. The plots demonstrate the rapid convergence of the MDPO-based algorithms.
@@
Test performance
Figure 1 and  2 show the test performance after 20,000 time steps depending on the distortion level. MDPO-Robust and PPO achieve the highest performance on the reward. However, since they do not optimise the Lagrangian, they perform poorly on the constraint-cost. MDPO-Robust-Augmented-Lag achieves the best overall performance by providing the lowest constraint-cost by far across all distortion levels. In the test performance after 200,000 time steps (see Figure 13 and  14 of Appendix F.1), the MCPMD algorithms are able to get to similar performance levels, and all three baselearners are competitive.
To summarise the overall performance quantitatively, Table 2(a) shows that MDPO-Robust is the top performer on the return after 20,000 time steps. MDPO-Robust-Lag and MDPO-Lag algorithms obtain similar levels of performance on the mean penalised return statistics, although the former obtain an improved minimum score, and both are far ahead of PPO and MCPMD algorithms. In Table 2(b), the performance after 200,000 time steps can be observed. Although MDPO-Robust-Lag is the top performer on the mean penalised return metrics, the algorithms now are much closer in test performance, and MCPMD-Lag and RMCPMD-Lag perform highest on the minimum performance. In summary, the solutions converge to relatively similar test performance levels but MDPO-based algorithms are superior in terms of sample efficiency.
@@
MDPO-Lag 37.3 $\pm$ 0.2 28.2 104.9 $\pm$ 6.8 23.8 35.5 $\pm$ 1.7 8.6
MDPO-Augmented-Lag 36.8 $\pm$ 0.2 27.4 104.9 $\pm$ 6.3 32.3 35.0 $\pm$ 1.6 13.8
@@
MDPO-Robust-Lag 34.7$\pm$ 0.3 25.3 106.9$\pm$ 4.7 57.9 34.3$\pm$ 1.1 25.3
MDPO-Robust-Augmented-Lag 34.2$\pm$ 0.3 24.0 107.6$\pm$ 4.3 68.1 33.8$\pm$ 1.2 24.0
@@
MDPO-Lag 42.5 $\pm$ 0.2 35.2 96.5 $\pm$ 9.2 -8.1 36.1 $\pm$ 3.9 -17.8
MDPO-Augmented-Lag 43.1 $\pm$ 0.2 36.0 95.5 $\pm$ 10.2 -19.9 35.0 $\pm$ 4.6 -28.3
@@
MDPO-Robust-Lag 39.5 $\pm$ 0.2 29.6 109.0 $\pm$ 6.7 39.0 37.3 $\pm$ 1.9 12.3
MDPO-Robust-Augmented-Lag 38.9 $\pm$ 0.2 29.1 106.8 $\pm$ 6.4 42.0 36.5 $\pm$ 1.8 13.1
@@ 5.2 Inventory Management
A second domain is the Inventory Management domain from Wang and Petrik, (2024), which involves maintaining an inventory of resources. The resources induce a period-wise cost and the goal is to purchase and sell resources such that minimal cost is incurred over time. To form a constrained variant of this benchmark, we introduce the constraint that the discounted sum of actions should not average to higher than zero (i.e. the rate of selling should not exceed that of purchasing). We use the same radial features and clipped uncertainty set as in the implementation on github https://github.com/JerrisonWang/JMLR-DRPMD. The implementation uses Gaussian parametric transition dynamics
@@
where $\mu_{\zeta_{1}}=(-4,5)$, $\mu_{\zeta_{2}}=(-2,8)$, and the uncertainty set is given by $\{\eta:\left\lVert\eta-\eta_{c}\right\rVert_{\infty}\leq\kappa\}$ where $\eta_{c}=(-2,3.5)$. With the exception of the constraint, and the number of steps per episode for the adversary, the setting matches Wang and Petrik, (2024). The constraint at time $t$ is given by
@@
The training performance can be found in Figure 9 of Appendix E.2 for non-robust algorithms and in Figure 10 of Appendix E.2 for robust algorithms. The plots confirm the effectiveness of the constrained optimisation algorithms and the robust algorithms can also find a good solution despite the environment being adversarial. It is also clear that the MDPO based algorithms converge more rapidly to solutions of the highest quality in constrained and robust-constrained optimisation, although there appears to be a small benefit of PPO in non-robust unconstrained optimisation.
@@
Test performance
Figure 5 and Figure 6 visualise the results for the test performance after 16,000 time steps. MDPO-Robust obtains the highest performance on the test reward, followed closely by MDPO and PPO. The constraint is challenging to satisfy for all algorithms, but MDPO-Lag, MDPO-Robust-Lag, and MDPO-Robust-Augmented-Lag obtain the best solutions with cost between 2 and 3 across the distortion levels. Algorithms based on PPO-Lag and MCPMD-Lag perform poorly with worst starting points and stronger performance degrading across distortion levels. After 400,000 time steps, the differences between the constrained algorithms is minimal in the test cases (see Figure 15 and  16 of Appendix F.2)
As can be observed in Table 3(a), MDPO-Robust, MDPO, and PPO are the top performers on the return after 16,000 time steps. All other statistics, including the mean and minimum of the signed and positive penalised return are maximised by MDPO-Robust-Augmented-Lag. In Table 3(b), the performance after the full 400,000 time steps can be observed. Although MDPO-Robust-Lag is the top performer in the penalised return metrics, the algorithms now are much closer in test performance. In summary, the solutions converge to relatively similar test performance levels but MDPO-based algorithms are superior in terms of sample efficiency.
@@
MDPO-Lag 68.5 $\pm$ 7.1 -162.8 -941.2 $\pm$ 33.6 -1106.8 -941.9 $\pm$ 32.7 -1106.8
MDPO-Augmented-Lag 75.2 $\pm$ 8.9 -184.6 -1366.6 $\pm$ 46.2 -1593.9 -1368.5 $\pm$ 43.8 -1593.9
RMCPMD-Lag 50.3 $\pm$ 27.4 -417.4 -2635.6 $\pm$ 113.0 -3159.5 -2638.3 $\pm$ $1$3159.5
@@
MDPO-Robust-Lag 60.5 $\pm$ 9.7 -224.5 -1081.1 $\pm$ 52.0 -1337.8 -1082.5 $\pm$ 50.7 -1337.8
MDPO-Robust-Augmented-Lag 66.9 $\pm$ 7.1 -162.5 -837.4 $\pm$ 21.2 -939.7 -838.1 $\pm$ 20.3 -939.7
@@
MDPO-Lag 66.9 $\pm$ 7.0 -159.6 -837.5 $\pm$ 21.1 -939.1 -838.1 $\pm$ 20.1 -939.1
MDPO-Augmented-Lag 66.8 $\pm$ 7.1 -167.2 -837.4 $\pm$ 21.4 -938.8 -838.0 $\pm$ 20.5 -938.8
@@
MDPO-Robust-Lag 60.9 $\pm$ 10.6 -221.8 -931.2 $\pm$ 32.0 -1117.7 -932.5 $\pm$ 30.6 -1117.7
MDPO-Robust-Augmented-Lag 66.9 $\pm$ 7.1 -164.5 -837.4 $\pm$ 21.3 -939.7 -838.0 $\pm$ 20.5 -939.7
@@
The training performance can be found in Figure 11 and Figure 12 of Appendix E.3. It is clear that the MDPO-based algorithm have improved sample-efficiency and converge to the highest unconstrained and constrained performances.
@@
Figure 5 and Figure 6 visualise the results for the test performance after 40,000 time steps. While all algorithms are sensitive to the distortion level, it can be seen that the robust algorithms have are generally shifted up for the rewards, and robust-constrained algorithms have been shifted down for the constraint-costs, indicating the effectiveness of robustness training. The MDPO-Robust algorithm is superior in the return and the the MDPO-Robust-Lag algorithms are found to be superior in the constraint-cost. While it is challenging to satisfy the constraint with limited training, after 400,000 time all algorithms have further improved, but remarkably MDPO-Robust-Augmented-Lag can satisfy all the constraints for all but the highest distortion levels (see Figure 17 and Figure 18 of Appendix F.3).
@@
MDPO-Lag 43.1 $\pm$ 0.5 -24.0 -2218.4 $\pm$ 446.4 -8609.9 -5483.8 $\pm$ 232.2 -9297.0
MDPO-Augmented-Lag 43.1 $\pm$ 0.5 -23.8 -2247.8 $\pm$ 442.3 -8594.0 -5479.1 $\pm$ 230.3 -9305.3
@@
MDPO-Robust-Lag 40.8 $\pm$ 0.5 -26.2 -1112.1 $\pm$ 444.2 -7558.3 -5353.0 $\pm$ 216.9 -8926.7
MDPO-Robust-Augmented-Lag 42.0 $\pm$ 0.5 -25.0 -1725.5 $\pm$ 443.2 -8179.6 -5450.3 $\pm$ 220.0 -9136.5
@@
PPO-Robust 60.7 $\pm$ 0.4 -2.2 -10691.3 $\pm$ 441.5 -17197.9 -11365.5 $\pm$ 364.6 -17245.6
MDPO-Robust 95.7 $\pm$ 0.4 36.2 -27546.2 $\pm$ 544.9 -35650.4 -27679.3 $\pm$ 522.0 -35651.8
@@
MDPO-Lag 35.0 $\pm$ 0.5 -30.7 1604.0 $\pm$ 420.5 -4504.8 -1138.5 $\pm$ 234.8 -5469.1
MDPO-Augmented-Lag 24.9 $\pm$ 0.5 -44.6 6318.8 $\pm$ 367.6 1555.7 -1146.3 $\pm$ 186.2 -4663.6
@@
MDPO-Robust-Lag 35.9 $\pm$ 0.4 -29.1 1100.4 $\pm$ 358.3 -3951.0 -2198.4 $\pm$ 198.4 -5525.9
MDPO-Robust-Augmented-Lag 26.0 $\pm$ 0.4 -39.1 5683.9 $\pm$ 346.5 902.2 -928.3 $\pm$ 128.2 -3117.7
The return and penalised return statistics after 40,000 time steps can be observed in Table 4(a). MDPO-Robust has the highest score on the mean and minimum of the test return, which indicates the robustness training was succesful in guaranteeing high levels of performance across the uncertainty set. MDPO-Robust-Lag outperforms all other algorithms on the penalised return statistics, with a large effect size on the signed penalised return. The other MDPO-Lag based algorithms (MDPO-Robust-Lag, MDPO-Augmented-Lag, MDPO-Lag) are close in performance on the (positive) penalised return. After 400,000 time steps, MDPO-Robust-Augmented-Lag has by far the highest mean and minimum positive penalised return while MDPO-Augmented-Lag has by far the highest mean and minmum signed penalised return (see Table 4(b). Overall, the data indicate the effectiveness for MDPO-based algorithms, and particularly the effectiveness of MDPO-Robust-Augmented-Lag for robust constrained optimisation.

### v2

@@ 6.1 Cartpole
To assess our algorithm on an RCMDP, we introduce the robust constrained variant of the well-known Cartpole problem (Barto et al.,, 1983; Brockman et al.,, 2016) by modifying the RMDP from Wang et al., (2024). The problem involves a mechanical simulation of a frictionless cart moving back and forth to maintain a pole, which is attached to the cart as an un-actuated joint, in the upright position. The agent observes $x,\dot{x},\theta,\dot{\theta}$ and takes actions in $\{\text{left},\text{right}\}$, which apply a small force moving either left or right. The agent receives a reward of +1+1 until either the cart goes out of the [2.4,2.4][-2.4,2.4]m bounds or the pole has an angle outside the range $12^{\circ},12^{\circ}]$. In contrast to the traditional problem, the mechanics are not deterministic, and following Wang et al., (2024), transitions dynamics models are multi-variate Gaussians of the form
@@
where $\mu_{c}(s,a)$ is the deterministic next state given $(s,a)\in\mathcal{S}\times\mathcal{A}$ based on the mechanics of the original Cartpole problem. We first train and evaluate algorithms in a CMDP problem for their ability to adhere to safety constraints. We then also show the benefits of robustness training and evaluate algorithms in an RCMDP problem.
@@
To incorporate robustness into our setting, we introduce delta-perturbations as in Wang et al., (2024), using an uncertainty set with parametrisation
@@
where $\delta\in[-\kappa_{i},\kappa_{i}]_{i=1}^{n}$. Our parameter settings differ in that we use a larger uncertainty set, with $\kappa_{i}$ being increased fivefold (making the problem more challenging) while the number of time steps is reduced to make the constraint satisfiable. The robust algorithms are trained on this uncertainty set, where the transition dynamics are adjusted based on the transition mirror ascent algorithm of Wang et al., (2024), i.e. by using Monte Carlo and mirror ascent. In the experiment, this amounts to a projected update, where the projection restricts the update to lie within the uncertainty set.
@@
The training performance of non-robust algorithms can be found in Appendix F.1 (Figure 7). The training performance of robust methods, which use LTMA, can be found in Figure 8. A trade-off in reward vs constraint-cost can be observed; that is, the constrained algorithms put emphasis on reducing the constraint-cost while the unconstrained only maximise the reward. The plots demonstrate the rapid convergence of the MDPO-based algorithms.
@@
Test performance
Figure 1 and  2 show the test performance after 20,000 time steps depending on the distortion level. MDPO-Robust and PPO achieve the highest performance on the reward. However, since they do not optimise the Lagrangian, they perform poorly on the constraint-cost. MDPO-Robust-Augmented-Lag achieves the best overall performance by providing the lowest constraint-cost by far across all distortion levels. In the test performance after 200,000 time steps (see Figure 13 and  14 of Appendix G.1), the MCPMD algorithms are able to get to similar performance levels, and all three baselearners are competitive.
@@
To summarise the overall performance quantitatively, Table 2(a) shows that MDPO-Robust is the top performer on the return after 20,000 time steps. MDPO-Robust-Lag and MDPO-Lag algorithms obtain similar levels of performance on the mean penalised return statistics, although the former obtain an improved minimum score, and both are far ahead of PPO-Lag and MCPMD-Lag algorithms. In Table 2(b), the performance after 200,000 time steps can be observed. MDPO-Robust-Lag, with or without augmentation, is the top performer on the mean penalised return metrics, followed by non-robust MDPO-Lag. MCPMD-Lag and RMCPMD-Lag perform highest on the minimum performance, and are followed by MDPO-Robust-Lag algorithms. In summary, the solutions converge to relatively similar test performance levels but MDPO-based algorithms are superior in terms of sample efficiency.
@@
MDPO-Lag 37.0 $\pm$ 0.2 28.0 106.1 $\pm$ 6.9 23.2 35.7 $\pm$ 1.6 11.9
MDPO-Augmented-Lag 37.5 $\pm$ 0.2 28.3 105.6 $\pm$ 7.2 18.4 35.9 $\pm$ 1.7 7.2
@@
MDPO-Robust-Lag 34.4 $\pm$ 0.3 24.9 108.4 $\pm$ 4.4 65.9 34.2 $\pm$ 1.2 24.9
MDPO-Robust-Augmented-Lag 34.6 $\pm$ 0.3 24.7 110.8 $\pm$ 4.9 61.6 34.6 $\pm$ 1.2 24.7
@@
MDPO-Lag 44.5 $\pm$ 0.2 37.8 90.9 $\pm$ 11.7 -38.9 36.0 $\pm$ 5.5 -42.4
MDPO-Augmented-Lag 45.2 $\pm$ 0.2 38.4 86.0 $\pm$ 12.5 -54.4 35.1 $\pm$ 6.6 -56.3
@@
MDPO-Robust-Lag 41.6 $\pm$ 0.2 32.1 105.1 $\pm$ 9.5 9.6 38.6 $\pm$ 2.5 2.6
MDPO-Robust-Augmented-Lag 41.2 $\pm$ 0.2 31.4 106.2 $\pm$ 9.1 13.7 38.6 $\pm$ 2.2 5.0
@@ 6.2 Inventory Management
A second domain is the Inventory Management domain from Wang et al., (2024), which involves maintaining an inventory of resources. The resources induce a period-wise cost and the goal is to purchase and sell resources such that minimal cost is incurred over time. To form a constrained variant of this benchmark, we introduce the constraint that the discounted sum of actions should not average to higher than zero (i.e. the rate of selling should not exceed that of purchasing). We use the same radial features and clipped uncertainty set as in the implementation on github https://github.com/JerrisonWang/JMLR-DRPMD. The implementation uses Gaussian parametric transition dynamics
@@
where $\mu_{\zeta_{1}}=(-4,5)$, $\mu_{\zeta_{2}}=(-2,8)$. The uncertainty set is given by $\{\eta:\left\lVert\eta-\eta_{c}\right\rVert_{\infty}\leq\kappa\}$ where $\eta_{c}=(-2,3.5)$. With the exception of the constraint, and the number of steps per episode for the adversary, the setting matches Wang et al., (2024). The constraint at time $t$ is given by
@@
The training performance can be found in Figure 9 of Appendix F.2 for non-robust algorithms and in Figure 10 of Appendix F.2 for robust algorithms. The plots confirm the effectiveness of the constrained optimisation algorithms and the robust algorithms can also find a good solution despite the environment being adversarial. It is also clear that the MDPO based algorithms converge more rapidly to solutions of the highest quality in constrained and robust-constrained optimisation, although there appears to be a small benefit of PPO in non-robust unconstrained optimisation.
@@
Test performance
Figure 5 and Figure 6 visualise the results for the test performance after 16,000 time steps. MDPO-Robust obtains the highest performance on the test reward, followed closely by MDPO and PPO. The constraint is challenging to satisfy for all algorithms, but MDPO-Lag based algorithms obtain the best solutions with cost between 2 and 3 across the distortion levels. Algorithms based on PPO-Lag and MCPMD-Lag perform poorly with worst starting points and stronger performance degrading across distortion levels. After 400,000 time steps, the differences between the algorithms is relatively small in the test cases (see Figure 15 and  16 of Appendix G.2)
@@
As can be observed in Table 3(a), MDPO-Robust is the top performer on the return after 16,000 time steps. All other statistics, including the mean and minimum of the signed and positive penalised return are maximised by MDPO-Lag based algorithms. In Table 3(b), the performance after the full 400,000 time steps can be observed. With the exception of PPO-Robust-Lag, all constrained algorithms converge to a similar optimum, indicating that the nominal environment and the other environments in the uncertainty set have large overlap. Similarly, the unconstrained algorithms reach a similar optimum. In summary, the solutions converge to relatively similar test performance levels but MDPO-based algorithms are superior in terms of sample efficiency.
@@
MDPO-Lag 66.8 $\pm$ 7.1 -167.2 -837.3 $\pm$ 21.4 -939.0 -837.9 $\pm$ 20.4 -939.0
MDPO-Augmented-Lag 66.9 $\pm$ 7.0 -159.4 -837.5 $\pm$ 21.1 -938.9 -838.0 $\pm$ 20.4 -938.9
RMCPMD-Lag 50.3 $\pm$ 27.4 -417.4 -2635.6 $\pm$ 113.0 -3159.5 -2638.3 $\pm$ 111.1 -3159.5
@@
MDPO-Robust-Lag 60.7 $\pm$ 10.7 -228.5 -930.9 $\pm$ 32.3 -1118.2 -932.2 $\pm$ 30.8 -1118.2
MDPO-Robust-Augmented-Lag 60.9 $\pm$ 10.6 -221.4 -938.6 $\pm$ 34.4 -1172.2 -940.1 $\pm$ 32.9 -1172.2
@@
MDPO-Lag 66.8 $\pm$ 7.2 -168.5 -837.3 $\pm$ 21.5 -939.3 -838.0 $\pm$ 20.4 -939.3
MDPO-Augmented-Lag 66.9 $\pm$ 7.1 -162.5 -837.5 $\pm$ 21.2 -939.2 -838.0 $\pm$ 20.4 -939.2
@@
MDPO-Robust-Lag 66.9 $\pm$ 7.1 -166.5 -837.4 $\pm$ 21.4 -939.3 -838.0 $\pm$ 20.4 -939.3
MDPO-Robust-Augmented-Lag 66.8 $\pm$ 7.1 -167.7 -837.2 $\pm$ 21.5 -939.1 -837.9 $\pm$ 20.5 -939.1
@@
The training performance can be found in Figure 11 and Figure 12 of Appendix F.3. It is clear that the MDPO-based algorithm have improved sample-efficiency and converge to the highest unconstrained and constrained performances.
@@
Figure 5 and Figure 6 visualise the results for the test performance after 40,000 time steps. While all algorithms are sensitive to the distortion level, it can be seen that the robust algorithms have are generally shifted up for the rewards, and robust-constrained algorithms have been shifted down for the constraint-costs, indicating the effectiveness of robustness training. The MDPO-Robust algorithm is superior in the return and the the MDPO-Robust-Lag algorithms are found to be superior in the constraint-cost. While it is challenging to satisfy the constraint with limited training, after 400,000 time all algorithms have further improved, but remarkably MDPO-Robust-Augmented-Lag can satisfy all the constraints for all but the highest distortion levels (see Figure 17 and Figure 18 of Appendix G.3).
@@
MDPO-Lag 43.4 $\pm$ 0.5 -24.6 -2391.4 $\pm$ 447.4 -8762.9 -5530.5 $\pm$ 237.8 -9481.2
MDPO-Augmented-Lag 42.8 $\pm$ 0.5 -23.5 -2126.5 $\pm$ 446.1 -8385.7 -5460.9 $\pm$ 229.5 -9142.2
@@
MDPO-Robust-Lag 41.4 $\pm$ 0.5 -24.6 -1416.9 $\pm$ 446.1 -7726.9 -5088.9 $\pm$ 221.7 -8560.6
MDPO-Robust-Augmented-Lag 39.5 $\pm$ 0.5 -29.6 -525.0 $\pm$ 441.2 -6861.9 -4957.0 $\pm$ 211.2 -8289.3
@@
PPO-Robust 60.9 $\pm$ 0.4 -2.4 -10802.7 $\pm$ 443.3 -17484.1 -11429.2 $\pm$ 367.4 -17506.1
MDPO-Robust 87.0 $\pm$ 0.4 24.4 -23312.6 $\pm$ 525.6 -31314.1 -23459.7 $\pm$ 499.3 -31314.1
@@
MDPO-Lag 34.9 $\pm$ 0.5 -32.5 1566.5 $\pm$ 420.9 -4406.7 -1180.5 $\pm$ 229.6 -5420.4
MDPO-Augmented-Lag 24.6 $\pm$ 0.5 -44.1 6421.1 $\pm$ 369.2 1517.0 -1139.5 $\pm$ 175.7 -4564.8
@@
MDPO-Robust-Lag 32.8 $\pm$ 0.5 -32.4 2617.9 $\pm$ 378.6 -2601.7 -1011.4 $\pm$ 169.9 -3767.4
MDPO-Robust-Augmented-Lag 26.8 $\pm$ 0.4 -38.6 5312.6 $\pm$ 340.2 758.7 -1219.2 $\pm$ 150.4 -3867.7
The return and penalised return statistics after 40,000 time steps can be observed in Table 4(a). MDPO-Robust has the highest score on the mean and minimum of the test return, which indicates the robustness training was succesful in guaranteeing high levels of performance across the uncertainty set. MDPO-Robust-Augmented-Lag outperforms all other algorithms on the signed and positive penalised return statistics. MDPO-Robust-Lag follows closely in mean and minimum performance on the positive penalised return. After 400,000 time steps, MDPO-Robust-Lag algorithms have by far the highest minimum penalised return while MDPO algorithms with augmented Lagrangian score remarkably well on the signed penalised return (see Table 4(b)). Overall, the data indicate the effectiveness for MDPO-based algorithms, and particularly the effectiveness of MDPO-Robust-Lag algorithms for robust constrained optimisation.